# LeTO: Learning Constrained Visuomotor Policy with Differentiable Trajectory Optimization

### v2

@@ III-D Policy Deployment
When the policy is deployed on the robot, the RNN is unrolled one-step at a time, $e_{t},h_{t+1}=E_{\theta}\left(o_{t},h_{t}\right)$. The hidden state $h$ is refreshed every $T_{s}$ steps. In addition, the training is on discrete trajectory clips while the robot trajectory is continuous in reality. Therefore, the forward pass of the DTO layer need to add a new constraint during test time to fully constrain the motion generation.
@@
where $\hat{v}_{t-1}$ represents the last executed action for the robot at time step $t$ (executed at time step $t-1$). Adding this new constraint is transferring the model trained on a discrete set of trajectories into continuous online trajectory optimization.
TABLE I: Summary of success rates. We use the imitation learning benchmark, RoboMimic (Visual Policy) [12]. We present the average of the maximum success rate (%) across 3 training seeds and 25 different initial environmental conditions (totaling 75 conditions). Results for IBC were sourced from [2] to enable a comparison with conventional methods. The black bold font indicates the highest success rate among all tasks. Notably, for the square task, LeTO achieves a comparable success rate to diffusion policy (bold in green) and surpasses diffusion policy with constraints clipping.
@@
LeTO $\textbf{100.0}\pm\textbf{ 0.0}$ $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}94.7}}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}\pm}\textbf{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}4.6}}$ $\textbf{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}88.0}}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}\pm}\textbf{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}4.0}}$
@@ IV Simulation Evaluation
We employed two simulation tasks used in [12] for our simulation evaluation: pick and place the can (can task) and grasping and assembling the square nut (square task), as shown in Fig. 4. The results demonstrate that our approach not only achieves success rates comparable to diffusion policy [2], but also significantly surpasses IBC [10] and LSTM+GMM [12]. Furthermore, owing to the presence of constraints and differentiable trajectory optimization, our method exhibits superior performance in terms of generated trajectory quality compared to other methods, which is very important for robot systems.
@@
The can and square tasks are from RoboMimic benchmark [12]. Apart from the most basic lifting task, we opted not to include the tool hanging and transport tasks. As highlighted in [40], solving optimization problems precisely, as we are doing here, exhibits cubic complexity concerning the number of variables and/or constraints. Consequently, our method’s training speed for tasks like tool hanging and transport, which involve large datasets or high dimension outputs, is sluggish, and under limited GPU computational resources, obtaining results within a short time is unfeasible. Nevertheless, as we have already demonstrated the capability of our approach to achieve high successful rates and generate high quality trajectories through the outcomes of the two tasks mentioned above, we decided against further training on the tool hanging and transport tasks.
We train policies on 3 seeds (42, 142, 242), and test in 25 different initial environments for each seeds (totaling 75 conditions). For each task, we train policies on two different types of datasets: Proficient-Human (ph), and Multi-Human (mh) datasets [12]. We report the success rate (see Table I) of all these experiments. We also report metrics for evaluating the quality of the generated trajectories (see Tables II, III, IV, and V).
TABLE II: Trajectory metrics ($10^{-2}$) of the can-ph task. The results correspond to the results in Table I.
@@
TABLE III: Trajectory metrics ($10^{-2}$) of the can-mh task. The results correspond to the results in Table I.
@@
TABLE IV: Trajectory metrics ($10^{-2}$) of the square-ph task. The results correspond to the results in Table I.
@@
TABLE V: Trajectory metrics ($10^{-2}$) of the square-mh task. The results correspond to the results in Table I.
@@
The velocity action outputs in RoboMimic dataset [12] are normalized to [-1, 1]. Therefore, for our implementation of LeTO, we impose a normalized velocity constraint [-1,1], and we introduce an acceleration constraint [-0.1,0.1]. For the simulation tests, we do not apply any position constraints, as our aim is to solely compare the success rates and the smoothness of the trajectories generated by different methods.
@@
where $i$ represents the $i$-th environment and $a$ represents acceleration. “lin” here represents linear motion and “rot” represents rotational motion. The remaining matrics can be computed by analogy. Additionally, it is worth noting that the values of the acceleration here are obtained by computing the difference of the normalized velocity. It is directly proportional to the actual acceleration, and this proportionality constant is a fixed value for all experiments. We mark the values within the range of 100% to 110% of the minimum value in each metric as the optimal indicators (bolded), and denote the number of optimal indicators for each method as “num-of-opt”.
@@
In order to investigate the impact of direct action clipping on success rate and trajectory smoothness for baselines, we implement a clipping method. It involves evaluating the difference between the generated action and the previously executed action to determine if it falls within the range of -0.1 to 0.1, which is the same constraints we implement for LeTO. If it exceeds this range, we apply clipping to bring it within this range. This approach constrains acceleration.
@@
We observe that LeTO significantly enhances trajectory smoothness by trajectory optimization while maintaining a consistently high success rate across all simulation tasks. LeTO achieves a success rate comparable to diffusion policy and surpassing LSTM-GMM and IBC. Interestingly, when applying clipping to constrain the actions generated by the diffusion policy and LSTM-GMM, although this method sometimes increases the smoothness of the trajectory, it results in a reduction in their task success rates.
@@
An intuitive explanation for this phenomenon is that both of these methods do not consider constraints during their training process; they merely fitted human demonstrations. When constraints are imposed on the actions generated by the trained models during deployment, it introduces additional factors that inevitably affect their performance. In contrast, our proposed LeTO, due to its ability to achieve end-to-end trajectory optimization while fitting demonstration data, balance the training objectives of satisfying constraints, smoothing the trajectories, and minimizing errors with demonstrations.
@@ V Realworld Evaluation
For real robot experiments, we tackled constraints-critical tasks that demand smooth and safe trajectories, which are Move-the-stack (Fig. 5) and Arrange-chopsticks (Fig. 7). We use SpaceMouse teleoperating the end-effector of the robot manipulator to collect human demonstration data. More details of experiments and results can be seen in the attached video.
@@ V-A Move-the-stack
TABLE VI: Metrics for evaluating the Move-the-stack task. “acc-peak” represents the highest acceleration achieved in any of the test trials while “avg-acc-mean/max/std” are the average of mean/max/std values across all test trials.
@@
[…]
TABLE VII: Metrics for evaluating the Arrange-chopsticks task. “contact-peak” represents the highest contact force achieved in any of the test trials while “avg-contact-max” is the average of max contact forces across all test trials.
@@
The objective of the “Move-the-stack” task is that the robot grasps a set of stacked objects, smoothly transport them, and ultimately place them onto a black board, ensuring none of the stacked objects fall off. The initial position of the robot and the black board are randomized. The policy inputs end-effector positions and images from a third-person perspective camera, and controls the end-effector position. We introduce velocity constraint [-1,1] and acceleration constraint [-0.25,0.25] for LeTO. Notably, while we did not place chopsticks on top during human demonstration collection, a chopstick was added atop the stacked cups during policy rollout.
For imitation learning algorithms, the task is challenging because: 1) The policy needs to generate very smooth trajectories. Any roughness in the trajectory can lead to vibrations at the robot’s end-effector, causing the cups to shake. This vibration can accumulate and lead to the chopstick dropping. 2) Since a chopstick is introduced during policy rollout, this out-of-distribution (OOD) observation adds greater uncertainty to the network’s input. This uncertainty may make the network’s output jerkier, resulting in a poorer quality trajectory.
The results are shown in Table VI. From the results, LeTO excels in terms of trajectory smoothness and, correspondingly, also boasts the highest success rate. This demonstrates that due to the presence of end-to-end trajectory optimization, LeTO can robustly maintain its effectiveness even when facing OOD data. Another point worth noting is that when chopsticks fall (which is also a type of OOD data), both the diffusion policy and LeTO can still move the stacked cups and place them on the black board. In contrast, LSTM-GMM often exhibits erroneous behaviors, such as moving to incorrect locations or acting erratically (see Fig. 6 and the attached video). This showcases the robustness of LeTO and the diffusion policy.

### v3

@@ III-D Policy Deployment
When the policy is deployed on the robot, the RNN is unrolled one-step at a time, $e_{t},h_{t+1}=E_{\theta}\left(o_{t},h_{t}\right)$. The hidden state $h$ is refreshed every $T_{s}$ steps. In addition, the training is done on discrete trajectory clips while the robot trajectory is continuous in reality. Therefore, the forward pass of the DTO layer need to add a new constraint during test time to fully constrain the motion generation.
@@
where $\hat{v}_{t-1}$ represents the last executed action for the robot at time step $t$ (executed at time step $t-1$). Adding this new constraint is to transfer the model trained on a discrete set of trajectories to continuous online trajectory optimization.
TABLE I: Experimental task setups and configurations. Initial States: Is the initial configuration (position and orientation) of the manipulated object randomized? Objective: Is the target object in the manipulation task randomized? For example, in the Square task, the stick to insert the nut remains stationary. PH and MH: The number of demonstrations in the dataset. PH represents proficient-human demonstrations, while MH stands for multi-human demonstrations. Our real-world task experiments were conducted exclusively on the PH dataset, consistent with the setup described in [15]. Act. Dim.: Action dimensions. HiPrec: Is high-precision manipulation required? Multi-Step: Does the task involve multiple sub-tasks? Num. of Cam.: The number of cameras used.
Task Initial States Objective PH MH Act. Dim. HiPrec Multi-Steps Num. of Cam.
Can Random Fixed 200 300 7 No No 2
Square Random Fixed 200 300 7 Yes No 2
Move-the-stack Random Random 100 - 3 Yes No 1
Arrange-chopsticks Random Fixed 120 - 4 No Yes 2
TABLE II: Training configuration of LeTO in simulation benchmarks. To improve learning efficiency, actions processed through LeTO are first normalized to a range between -1 and 1 [29]. Consequently, the velocity and acceleration constraints applied here pertain to the normalized actions.
Velocity Constraints [${v}_{\text{min}},{v}_{\text{max}}$] Acceleration Constraints [${a}_{\text{min}},{a}_{\text{max}}$] Smoothing Weight $\alpha$ Sampling Length $T_{s}$ Predicted Length $T_{p}$
[-1,1] (normalized values) [-0.1,0.1] (normalized values) 1 12 6
TABLE III: Summary of success rates. We use the imitation learning benchmark, RoboMimic (Visual Policy) [29]. We present the average of the maximum success rate (%) across 3 training seeds and 25 different initial environmental conditions (totaling 75 conditions). Results for IBC were sourced from [15] to enable a comparison with conventional methods. The black bold font indicates the highest success rate among all tasks. Notably, for the square task, LeTO achieves a comparable success rate to diffusion policy (bold in green) and surpasses diffusion policy with constraints clipping.
@@
LeTO $\textbf{100.0}\pm\textbf{ 0.0}$ $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}94.7}}{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}\pm}\textbf{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}4.6}}$ $\textbf{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}88.0}}{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}\pm}\textbf{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}4.0}}$
@@ IV Simulation Evaluation
IV-A Experimental Setup
We employed two simulation tasks used in [29] for our simulation evaluation: pick and place the can (can task) and grasping and assembling the square nut (square task), as shown in Fig. 4. More details of the task configurations can be seen in Table I. The results demonstrate that within these two simulation tasks, LeTO not only achieves success rates comparable to diffusion policy [15], but also significantly surpasses IBC [14] and LSTM+GMM [29]. Furthermore, owing to the presence of constraints and differentiable trajectory optimization, our method exhibits superior performance in terms of generated trajectory quality compared to other methods, which is very important for robot systems.
@@ IV-A Experimental Setup
The can and square tasks are from the RoboMimic benchmark [29]. Apart from the most basic lifting task, we opted not to include the tool hanging and transport tasks. As highlighted in [44], solving optimization problems precisely, as we are doing here, exhibits cubic complexity concerning the number of variables and/or constraints. Consequently, our method’s training speed for tasks like tool hanging and transport, which involve larger datasets or high-dimensional output, is sluggish, and under limited GPU computational resources, obtaining results within a short time is unfeasible. Nevertheless, as we have already demonstrated the capability of our approach to achieve high successful rates and generate high-quality trajectories through the outcomes of the two tasks mentioned above, we decided against further training on the tool hanging and transport tasks.
We train policies on 3 seeds (42, 142, 242) with batch size 64, and test in 25 different initial environments for each seed (totaling 75 conditions). For each task, we train policies on two different types of datasets: Proficient-Human (ph), and Multi-Human (mh) datasets [29]. We report the success rate (see Table III) of all these experiments. We also report metrics to evaluate the quality of the generated trajectories (see Tables IV, V, VI, and VII).
TABLE IV: Trajectory metrics ($10^{-2}$) of the can-ph task. The results correspond to the results in Table III. These values represent normalized accelerations. To convert the results back to their raw form, use a scalar of 20 $\text{m}/\text{s}^{2}$ for linear acceleration and 200 $\text{rad}/\text{s}^{2}$ for rotational acceleration. For example, the maximum linear acceleration for LSTM-GMM is $10^{-2}\times 77.49\times 20~{}\text{m}/\text{s}^{2}=15.50~{}\text{m}/\text{s}^{2}$ (which exceeds the gravitational acceleration), whereas for LeTO it is only $10^{-2}\times 10\times 20~{}\text{m}/\text{s}^{2}=2~{}\text{m}/\text{s}^{2}$.
@@
TABLE V: Trajectory metrics ($10^{-2}$) of the can-mh task. The results correspond to the results in Table III. These values represent normalized accelerations. To convert the results back to their raw form, use a scalar of 20 $\text{m}/\text{s}^{2}$ for linear acceleration and 200 $\text{rad}/\text{s}^{2}$ for rotational acceleration.
@@
TABLE VI: Trajectory metrics ($10^{-2}$) of the square-ph task. The results correspond to the results in Table III. These values represent normalized accelerations. To convert the results back to their raw form, use a scalar of 20 $\text{m}/\text{s}^{2}$ for linear acceleration and 200 $\text{rad}/\text{s}^{2}$ for rotational acceleration.
@@
TABLE VII: Trajectory metrics ($10^{-2}$) of the square-mh task. The results correspond to the results in Table III. These values represent normalized accelerations. To convert the results back to their raw form, use a scalar of 20 $\text{m}/\text{s}^{2}$ for linear acceleration and 200 $\text{rad}/\text{s}^{2}$ for rotational acceleration.
@@
TABLE VIII: The results of LeTO on square-ph with varying acceleration constraints. To enhance readability, we have highlighted the optimal values in each metric in bold red and the second-optimal values in bold green. The acceleration results employ the same normalization as outlined in Tables IV, V, VI, and VII
0.05 0.1 0.2 0.5 1
succ-rate 0.88 0.96 0.96 0.92 0.88
avg-mean-lin 0.033 0.049 0.061 0.071 0.067
avg-max-lin 0.05 0.10 0.20 0.46 0.58
avg-std-lin 0.017 0.033 0.051 0.071 0.072
avg-mean-rot 0.012 0.013 0.012 0.013 0.013
avg-max-rot 0.050 0.090 0.121 0.145 0.134
avg-std-rot 0.011 0.014 0.014 0.015 0.015
For the CNN+LSTM backbone of LeTO, we employ the hyperparameters selected in [29]. The other hyperparameters for training LeTO are detailed in Table II. To improve learning efficiency, actions processed through LeTO are first normalized to a range between -1 and 1 [29]. Consequently, the velocity and acceleration constraints applied here pertain to the normalized actions. See the codebase [^1] https://github.com/ARISE-Initiative/robosuite for more information of action unnormalization.
Setting velocity constraints within the range of [-1, 1] ensure that the trajectories generated by LeTO during both training and inference do not exceed the maximum velocity in the human demonstration data. This is a logical approach since we can assume that the maximum velocity in a high-quality human demonstration dataset is reasonable and does not lead to dangerous or unstable behavior in the robot.
The acceleration constraint of [-0.1, 0.1] effectively imposes additional smoothness on the trajectories. As acceleration is not directly controlled, the trajectories in the human demonstration may not be sufficiently smooth. In LeTO, imposing hard acceleration constraints optimizes the generation of higher-quality, smoother trajectories while ensuring manipulation performance.
For the simulation tests, we do not apply any position constraints, as our aim is to solely compare the success rates and the smoothness of the trajectories generated by different methods.
@@
where $i$ represents the $i$-th environment and $a$ represents acceleration. “lin” here represents linear motion and “rot” represents rotational motion. The remaining matrices can be computed by analogy. Additionally, it is worth noting that the acceleration values here are obtained by computing the difference of the normalized velocity. It is directly proportional to the actual acceleration, and this proportionality constant is a fixed value for all experiments. We mark the values within the range of 100% to 110% of the minimum value in each metric as the optimal indicators (bolded), and denote the number of optimal indicators for each method as “num-of-opt”.
@@
In order to investigate the impact of direct action clipping on success rate and trajectory smoothness for baselines, we implement a clipping method. It involves evaluating the difference between the generated action and the previously executed action to determine whether it falls within the range of -0.1 to 0.1, which are the same constraints that we implement for LeTO. If it exceeds this range, we apply clipping to bring it within this range. This approach restricts acceleration.
IV-B Results and Analysis
As shown in Tables IV, V, VI, and VII, it can be observed that LeTO significantly enhances trajectory smoothness by trajectory optimization. For example, if we unnormalize the values to true values for comparison, in Table IV, the maximum linear acceleration for LSTM-GMM is $10^{-2}\times 77.49\times 20~{}\text{m}/\text{s}^{2}=15.50~{}\text{m}/\text{s}^{2}$ (which exceeds the gravitational acceleration), while for LeTO it is only 2 $\text{m}/\text{s}^{2}$. The average acceleration of LSTM-GMM is also about twice that of LeTO. These metrics indicate that the trajectories generated by LSTM-GMM are significantly jerkier than those produced by LeTO. The same analysis applies to diffusion policy, and we can obtain similar results.
@@ IV-B Results and Analysis
Moreover, LeTO achieves a success rate comparable to that of diffusion policy and surpasses LSTM-GMM and IBC in these tasks. Interestingly, when clipping is applied to constrain the actions generated by the diffusion policy and LSTM-GMM, although this method sometimes increases the smoothness of the trajectory, it results in a reduction in their task success rates.
@@
An intuitive explanation for this phenomenon is that both methods do not consider constraints during their training process; they merely fitted human demonstrations. When constraints are imposed on the actions generated by the trained models during deployment, it introduces additional factors that inevitably affect their performance. In contrast, our proposed LeTO, due to its ability to achieve end-to-end trajectory optimization while fitting demonstration data, balances the training objectives of satisfying constraints, smoothing the trajectories, and minimizing errors with demonstrations.
While all the policies we benchmark might successfully complete these tasks in simulation, constraint guarantees are crucial metrics for robotic systems. People aim for robotic systems to operate safely and reliably over the long term, not just to complete a task once. Violating constraints and producing jerky trajectories could lead to unsafe situations because neural networks are inherently uninterpretable black boxes. Moreover, if a robot consistently executes jerky trajectories to complete tasks, it places a greater burden on the hardware. Over time, these factors undermine the robustness and stability of the entire system.
Historically, trajectory optimization for robots has focused on meeting constraints to enhance system robustness. In our paper, LeTO can be seen as a ”gray box” that offers a certain level of interpretability and can perform end-to-end trajectory optimization with hard constraints. This represents our attempt to connect trajectory optimization with robot learning. Moreover, we demonstrated that simply constraining the policy through clipping results in a decrease in performance. Therefore, we believe it is meaningful to explore how to integrate constraints into the policy to balance the objectives of skill learning and trajectory optimization with constraint guarantees. LeTO provides a practical example of how to achieve this integration.
To investigate the impact of acceleration constraints on the results of LeTO, we have conducted an ablation study on acceleration constraints (see Table VIII for details). Through this ablation study, we can the following observations.
1. Different acceleration constraints affect the success rate of the policy. Both overly restrictive and overly lenient constraints can reduce the success rate, resulting in a trend where the success rate initially increases with the constraint values before decreasing.
2. Table VIII demonstrates that LeTO, when implementing tight acceleration constraints, yields trajectories with notably superior smoothness based on differentiable optimization. The acceleration constraints determines the smoothness of the resulting trajectories. Tighter acceleration constraints result in smoother trajectories. As the maximum acceleration in the constraints approaches the maximum acceleration in the dataset, The trajectories generated by LeTO progressively approximate the original trajectories in the dataset.
3. Specifically, Tables  VI and VIII show that both the unconstrained LSTM-GMM and diffusion policy, and the loosely constrained LeTO ultimately generate a maximum linear acceleration of around 0.6 and a maximum rotational acceleration of about 0.14. Under relaxed or no constraints, the trajectories generated by the learned policies resemble those in the dataset across various metrics. This indicates that the dataset trajectories’ linear and rotational accelerations are approximately around 0.6 and 0.14, respectively (all values are normalized). As the constraints in LeTO transition from 1 to 0.05, the maximum linear acceleration of the trajectories it generates also transitions from 0.58 to 0.05, and the rotational acceleration transitions from 0.134 to 0.05.
4. This ablation study highlights LeTO’s superiority in terms of interpretability and controllability. When adjusting constraints, the metrics of the trajectories it generates remain consistent with the design of the differentiable trajectory optimization layer. Due to the black-box nature of neural networks, they lack the capability to “control” the final outputs, being limited only to fitting datasets. LeTO’s adjustable influence on trajectories stems from the theoretical fulfillment of constraints. This ensures that regardless of the input observations, the trajectories output by LeTO will always satisfy the given constraints by trajectory optimization. Such theoretical assurances enhance the interpretability of the entire model and are crucial for long-term safety and robustness.
@@ V Realworld Evaluation
For real robot experiments, we tackled constraints-critical tasks that demand smooth and safe trajectories, which are Move-the-stack (Fig. 5) and Arrange-chopsticks (Fig. 7). We use SpaceMouse [^2] https://3dconnexion.com/dk/product/spacemouse-compact/ teleoperating the end-effector of the robot manipulator to collect human demonstration data. SpaceMouse is widely used for data collection in imitation learning-related research [15, 29, 5, 46]. By manipulating the SpaceMouse, we can control the six degrees of freedom in the robot’s end-effector velocity, as well as the gripper’s grasping and releasing actions, thereby enabling the collection of data for various manipulation tasks. During the data collection process, the robot state information and images from the cameras are recorded simultaneously. More details of the task configurations can be seen in Table I. More details of experiments and results can be seen in the attached video.
@@ V-A Move-the-stack
[…]
The objective of the “Move-the-stack” task is that the robot grasps a set of stacked objects, smoothly transports them, and ultimately places them onto a black board, ensuring none of the stacked objects fall off. The initial positions of the robot and the black board are randomized. The policy inputs end-effector positions and images from a third-person perspective camera and controls the end-effector position. We introduce the velocity constraint [-1,1] and the acceleration constraint [-0.25,0.25] for LeTO. In particular, while we did not place chopsticks on top during human demonstration collection, a chopstick was added to the stacked cups during policy rollout.
For imitation learning algorithms, the task is challenging because: 1) The policy needs to generate very smooth trajectories. Any roughness in the trajectory can lead to vibrations at the robot’s end-effector, causing the cups to shake. This vibration can accumulate and cause the chopstick to drop. 2) Since a chopstick is introduced during policy rollout, this out-of-distribution (OOD) observation adds greater uncertainty to the network input. This uncertainty may make the network’s output jerkier, resulting in a poorer quality trajectory.
TABLE IX: Metrics for evaluating the Move-the-stack task. “acc-peak” represents the highest acceleration achieved in any of the test trials while “avg-acc-mean/max/std” are the average of mean/max/std values across all test trials.
@@
TABLE X: Metrics for evaluating the Arrange-chopsticks task. “contact-peak” represents the highest contact force achieved in any of the test trials while “avg-contact-max” is the average of max contact forces across all test trials.
@@
The results are shown in Table IX. From the results, LeTO excels in terms of trajectory smoothness and, correspondingly, also boasts the highest success rate. This shows that due to the presence of end-to-end trajectory optimization, LeTO can robustly maintain its effectiveness even when facing OOD data. Another point worth noting is that when chopsticks fall (which is also a type of OOD data), both the diffusion policy and LeTO can still move the stacked cups and place them on the black board. In contrast, LSTM-GMM often exhibits erroneous behaviors, such as moving to incorrect locations or acting erratically (see Fig. 6 and the attached video).
The Move-the-stack task highlights the importance of trajectory optimization. For methods solely based on neural networks, there is a lack of control over the generated trajectories, resulting in trajectories that are less smooth and more uncertain. The inherent uninterpretability of neural networks introduces greater uncertainty.